OpenStudy (anonymous):

Can you please help me get through this question? I really want to learn how to do this :(

OpenStudy (anonymous):

So the question is 3(x+1)= -2(x-1)+6

OpenStudy (anonymous):

3x+3=-2x+2+6 3x+2x=8-3 5x=5 x=5/5 x=1

OpenStudy (anonymous):

Wait how did the equation turn into 3x+3=-2x+2+6?

OpenStudy (anonymous):

multiply three into x-1....i mean open the brckets by multiplying

OpenStudy (anonymous):

3x+3 is the product of 3 * (x+1)

OpenStudy (anonymous):

yes, amirreza is right

OpenStudy (anonymous):

Okay so another thing

OpenStudy (anonymous):

Why does the negative 2 become positive?

OpenStudy (anonymous):

minus and minus multiply nad become plus

OpenStudy (anonymous):



OpenStudy (anonymous):